OpenStudy (anonymous):

solve the equation for x. radical 7- radical x+3=2

OpenStudy (anonymous):

\[(-\sqrt{x+3})^2=(2+\sqrt{7})^2\]

OpenStudy (anonymous):

x+3=11+4sqrt(7)...............\[x=8+4\sqrt(7)\]

OpenStudy (anonymous):

the question is \[\sqrt{7-\sqrt{x+3}}=2\]